1.) Bari jay a gown manufacturer recieved an order for prom dresses from china. Her cost is $35 per gown. if her markup based on selling price is 81% what is the selling price of each gown?

2.) Ben supple an importer in brownsville, has just received a shipment of peruvian opals that he is pricing for sale. He paid $104 for the shipment. Assume he wants a 75% markup. calculate the selling price when the markup is based on the selling price and also calculate the selling price when the markup is based on cost.
Answer #1

Selling price = Cost price + Mark up

Selling price = $35 + 81%*Selling price

Hence, Selling price = 35/19%

= $184.2105

i.e. $184.21

2)Cost = $104

Selling price when mark up is based on selling price = 104/(1-75%) = $416

Selling price when mark up is based on cost = 104 + 104*75% = $182